OpenStudy (31356):

x=y+8 in standard form? MEDAL REWARDED!!

OpenStudy (anonymous):

y = x-8

OpenStudy (smileyxl3):

x-y-8 = 0

OpenStudy (anonymous):

Standard form= Ax+By=C. You subtract the y from the latter side and that would give you x-y=8

OpenStudy (anonymous):

lalohao and smileyx13 are both wrong

OpenStudy (31356):

Thx DaBaller